OpenStudy (anonymous):

Determine the value of R: (A) m=4, (4, R), (-4, -7) (A) What is the slope? (B) m=3, (3, 1), (-3, R)

OpenStudy (ash2326):

a) given slope=4 we know slope = ( y2-y1)/(x2-x1) if we are given two points x2,y2 and x1,y1 slope m=4 4=(r-(-7))/(4-(-4)) 4=(r+7)/(8) 32=r+7 r=25 b)m= 3 (3,1) and (-3, R) 3=(R-1)/(-3-3) 3=(R-1)/(-6) -18+1=R R=-17